Question 3 (15 Marks): Consider a closed economy, where the price level P is constant and equal to 5.You are given the following additional information:

Cd = 1,000 + 0.8(Y - T)

T = 0.2Y – 100

Id = 2,000 – 1,000r

G = 600

L = 20,000r

M = 5,000

A. (5 Marks): What is the equilibrium level of income? What is the equilibrium level of the interest rate? Explain how you obtained your answers.

B. (5 Marks): The public increases its demand for money such that L = 30,000r. The Bank of Canada responds by increasing the money supply such that M = 7,500. What is the equilibrium level of income? What is the equilibrium level of the interest rate? Explain how you obtained your answers.

C. (5 Marks): Compare the equilibrium level of income for Part A with the equilibrium level of income for Part B. Explain why this result has occurred.
Answer #1

(PART A)

Goods market equilibrium condition: Y = C + Id + G

Y = 1,000 + 0.8[Y - (0.2Y - 100)] + 2,000 - 1,000r + 600

Y = 3,600 + 0.8(Y - 0.2Y + 100) - 1,000r

Y = 3,600 + 0.8(0.8Y + 100) - 1,000r

Y = 3,600 + 0.64Y + 80 - 1,000r

0.36Y = 3,680 - 1,000r

Y = (3,680 - 1,000r) / 0.36.........(equation of IS curve)

Money market equilibrium condition: Ld = Ms/P

20,000r = 5,000/5 = 1,000

r = 0.05

Substituting in IS equation, we get

Y = [3,680 - (1,000 x 0.05)] / 0.36 = (3,680 - 50) / 0.36 = 3,630 / 0.36 = 10,083

(PART B)

Revised money market equilibrium condition:

30,000r = 7,500/5 = 1,500

r = 0.05

Substituting in IS equation, we get

Y = [3,680 - (1,000 x 0.05)] / 0.36 = (3,680 - 50) / 0.36 = 3,630 / 0.36 = 10,083

(PART C)

Equilibrium interest rate and income are both same in (A) and (B).

Increase in interest rate due to an increase in demand for money is exactly equal to the decrease in interest rate due to an increase in money supply, hence as net effect, interest rate is the same. As a result, equilibrium output is unchanged.